OpenStudy (anonymous):

Which of the following would be the best first step to solve the system of equations? 2x – 9y = 6 4x – 3y = –7

OpenStudy (anonymous):

combine like terms ex.\[62x+6y=-7\]

OpenStudy (anonymous):

Pair of linear equations in 2 variables. 1. Find the value of one variable in terms of the other. 2x=6+9y x=(6+9y)/2 2. Substitute the obtained value in the second equation. 4(6+9y)/2-3y=-7 3. Find y and then, x.